OpenStudy (anonymous):

An auditorium earned $25,000 in sold-out concert ticket sales. Front section tickets cost $75 per seat and back section tickets cost $50 per seat. The number of front section seats is twice the number of back section seats. How many seats are in the front section?

OpenStudy (anonymous):

f = 2b ->75* (2b) + 50b = 25,000 => 150b + 50b = 25,000 => b = 25,000/ 200 = 125 back seats Thus f = 2 * 125 = 250 front seats